OpenStudy (anonymous):

Will give medal! Help with complex numbers! i√(5+3) / i√(5-3), write in the for a+bi

OpenStudy (solomonzelman):

1) Simplify the fractions \[\frac{i}{8} \div \frac{i}{2}~~~~~~->~~~~~~\frac{i}{8} \times\frac{2}{i}~~~~~~->~~~~~~\frac{2i}{8i}~~~~~~->~~~~~~\frac{1}{4}\]

OpenStudy (anonymous):

apparently the answer is